Sunlight Company needs a machine for its manufacturing process. The cost of the new machine is $80,700. The expected useful life of the machine is 8 years. At the end of 8-year period, the machine would have no salvage value. After installation, the machine would increase cash inflows by $30,000 per year. Sunlight is interested to know the net present value of the machine to accept or reject this investment. The minimum required MARR of the company is 16% on all capital investments. Compute net present value of the machine. Is it acceptable to purchase the machine?
Answer #1

ANSWER:

I = 16% AND N = 8 YEARS

PW = INITIAL COST + CASH INFLOWS PER YEAR(P/A,I,N)

PW = -80,700 + 30,000(P/A,16%,8)

PW = -80,700 + 30,000 * 4.34

PW = -80,700 + 130,307.73

PW = 49,607.73

SINCE THE PW IS POSITIVE AND SO IT IS ACCEPTABLE TO PURCHASE THE MACHINE.